Suppose an economy has no income taxes or imports. How is the size of the expenditure multiplier related to the marginal propensity to consume? What is the multiplier if the MPC equals 0.25? If the MPC equals 0.50? If the MPC equals 0.90?
Answer #1

ANswer

Multiplier =1/(1-MPC)

the multiplier increases as the MPC increases

MPC=0.25

Multiplier =1/(1-0.25)=1.33333333=1.33

MPC=0.5

Multiplier =1/(1-0.5)=2

MPC=0.9

Multiplier =1/(1-0.9)=10